21) Find the maximum value of the function f (z)=e^3z on the curve C:|z|=2.
Explain the problem with step by step process?
1
Expert's answer
2021-09-16T02:11:03-0400

Let"z_j =x_j +i_j," where "xj,yj" are real number=1,2,3.

 (a) We have

"z1 +z2 =(x1 +x2)+i(y1 +y2)=(x2 +x1)+i(y2 +y1)=z2 +z1." "z1z2 =(x1x2 \u2212y1y2)+i(x1y2 +x2y1)=(x2x1 \u2212y2y1)+i(y2x1 +y1x2)=z2z1."



is valid for real numbers and so it holds for the real and imaginary parts of"z_1z_2,z_3."

Consequently we have

"(z_1 +z_2)+z_3 =z_1 +(z_2 +z_3)."